OpenStudy (briana.img):

The social studies teacher is giving a test worth 80 points, and containing 34 questions. There are 2 point and 4 point questions on the test. How many of each type of question are on the test?

OpenStudy (briana.img):

i tried doing it but i ended up just confusing myself

OpenStudy (anonymous):

so you want to know how many 2 point questions and how many 4 point questions there are. That means that those are your two variables. Lets say the number of 2 point questions is represented by x and the number of 4 point questions is represented by y Now since we have 2 variables, we need to equations to solve for each. First, we know that the test is a total of 80 points. You get 2 points for every x question so we'll write that as 2*x and 4 points for every y question so 4*y and the total is 80 therefore: 2*x + 4*y = 80 We also know that there is a total of 34 questions. therefore: x + y = 34 Your system of 2 equations in 2 variables representing the problem is therefore: 2*x + 4*y = 80 x + y = 34 There are lots of ways of solving from here. Do you know any?

OpenStudy (briana.img):

@LifeEngineer method of substitution, method of elimination, and graphing?

OpenStudy (cwrw238):

i think elimination would do fine for this one

OpenStudy (cwrw238):

multiply the second equation by -2 and add

OpenStudy (briana.img):

x+y=34*-2 ?

OpenStudy (briana.img):

-68

OpenStudy (cwrw238):

2x + 4y = 80 -2x - 2y = -68

OpenStudy (cwrw238):

add ing these 2 equations will eliminate x

OpenStudy (briana.img):

@cwrw238 2x+4y=80+(-2)x-2y=-68?

OpenStudy (cwrw238):

yes

OpenStudy (cwrw238):

subtract them term by term to get an equation in y only

OpenStudy (briana.img):

@cwrw238 4x+2y=80+(-68)?

OpenStudy (cwrw238):

no